OpenStudy (anonymous):

Prove by induction: n ∑(i/2^i) = 2 - (n+2)/2^n i=1

OpenStudy (anonymous):

ah a repeat. did you start with proving true for n = 1?

OpenStudy (anonymous):

then you assume true for n, show true for n + 1 so \[\sum_{i=1}^{n+1}\frac{i}{2^i}=\sum_{i=1}^n\frac{i}{2^i}+\frac{n+1}{2^{n+1}}\]

OpenStudy (anonymous):

by induction the first term is \[2-\frac{n+2}{2^n}\] so you have \[\sum_{i=1}^{n+1}\frac{i}{2^i}=\sum_{i=1}^n\frac{i}{2^i}+\frac{n+1}{2^{n+1}}=2-\frac{n+2}{2^n}+\frac{n+1}{2^{n+1}}\]

OpenStudy (anonymous):

now compute \[\frac{n+1}{2^{n+1}}-\frac{n+2}{2^n}\] and you will get exactly what you want

OpenStudy (anonymous):

that's the part i'm having difficulty with

OpenStudy (anonymous):

namely \[ \frac{n+1}{2^{n+1}}-\frac{n+2}{2^n}=\frac{n+1-2(n+2)}{2^{n+1}}\]

OpenStudy (anonymous):

i multiplied top and bottom of second fraction by 2

OpenStudy (anonymous):

multiply out and get \[\frac{n+1-2n-4}{2^{n+1}}=\frac{-n-3}{2^{n+1}}=-\frac{n+3}{2^{n+1}}\]

OpenStudy (anonymous):

which is exactly what you needed to get to finish the induction